OpenStudy (anonymous):

Use an identity to solve each equation on interval [0,2pi) 2cos2x-sinx-1=0

sam (.sam.):

Is this your question?\[2\cos(2x)-sinx-1=0?\]

OpenStudy (anonymous):

yes

sam (.sam.):

Ok we have to solve this by factoring, and we need to change the cosine into sine to match it, so we can use the identity \(\cos(2x)=1-2\sin^2(x)\) \[2(1-2\sin^2(x))-\sin(x)-1=0 \\ \\ 2-4\sin^2(x)-\sin(x)-1=0 \\ \\ -4\sin^2(x)-\sin(x)+1=0\] We can factor it now, can you continue?

OpenStudy (anonymous):

No this is the part i suck at the mostt :(

sam (.sam.):

Oh when you can't factor it, use quadratic formula.

sam (.sam.):

\[-4\sin^2(x)-\sin(x)+1=0 \\ \\ Ax^2+Bx+C=0 \\ \\ \text{When} \\ \\ A=-4, B=-1, C=1\] \[\sin(x)=\frac{-B \pm \sqrt{B^2-4AC}}{2A}\]

OpenStudy (anonymous):

so it will be: \[\frac{ -1\pm \sqrt{-15} }{ -8 }\]

sam (.sam.):

Careful

sam (.sam.):

\[\sin(x)=\frac{-(-1) \pm \sqrt{(-1)^2-4(-4)(1)}}{2(-4)}\]

OpenStudy (anonymous):

Positive 1

sam (.sam.):

\[\sin(x)=\frac{1 \pm \sqrt{17}}{-8}\]

sam (.sam.):

Then, we have 2 sines out of this \[\sin(x)=-\frac{1}{8}+\frac{\sqrt{17}}{8}~~~~~~or~~~~~~\sin(x)=-\frac{1}{8}-\frac{\sqrt{17}}{8}\]

sam (.sam.):

Inverse that will give you x \[x=\sin^{-1}(-\frac{1}{8}+\frac{\sqrt{17}}{8})~~~~~or~~~~~x=\sin^{-1}(-\frac{1}{8}+\frac{\sqrt{17}}{8})\]

OpenStudy (anonymous):

So after i find x how will i solve for the intervals between [0,2pi)

sam (.sam.):

Use a calculator to find that, you should get x=22.97 and 39.8

sam (.sam.):



sam (.sam.):

For positive angle, look at the positive side x=180-22.97=157 For negative angle is the negative side x=180+39.8=219.8 x=320.2 So we got x=22.97,157,219.8,320.2 39.8 doesn't include because its negative

sam (.sam.):

and also, you have to convert the angles to radians

OpenStudy (anonymous):

In my homework they have it in radians. I should just convert?

sam (.sam.):

\[\frac{\theta}{180} \times \pi=randians\] \[\frac{radians}{\pi} \times 180=degrees\]